OpenStudy (samara89):

A cylindrical pillow has a diameter of 30 cm and a height of 40 cm. It is 80% filled with spherical foam pellets, each with a radius of 3 cm. Approximately how many pellets are in the pillow? Use 3.14 to approximate pi and express your answer to the nearest whole number.

OpenStudy (samara89):

@welshfella

OpenStudy (samara89):

is the answer 200 pellets

OpenStudy (greatlife44):

Calculate the total volume of the cylinder. \[\pi*r^{2}*h = Volume~cylinder \] Then multiply this by 0.8 ~ 80% to find the volume occupied by the pellets. \[\pi*r^{2}*h*0.8 = volume~occupied~by~pellets \] now calculate the volume a sphere you know the radius 3cm \[\frac{ 4 }{ 3 }\pi*r^{3} = volume~of~a~sphere \]

OpenStudy (samara89):

is the answer 200 yes or no that is my answer

OpenStudy (welshfella):

what do you get fo the volume of the cylinder * 0.8?

OpenStudy (greatlife44):

we will walk you through it but it's against OS code of conduct to give direct answers

OpenStudy (samara89):

do not know

OpenStudy (welshfella):

?? you must have worked that out?

OpenStudy (welshfella):

you had to in order to get an answer

OpenStudy (welshfella):

greatlife gave you the formula

OpenStudy (samara89):

22608

OpenStudy (welshfella):

yes tahts correct

OpenStudy (samara89):

22608/113.04=200

OpenStudy (welshfella):

now you need to divide that by the valume of one pellet , right?

OpenStudy (welshfella):

yes

OpenStudy (samara89):

so 200 is the answer

OpenStudy (welshfella):

it is . You should have more confidence in yourself.

OpenStudy (welshfella):

I think you are more able than you think you are.

OpenStudy (samara89):

I have more questions

OpenStudy (welshfella):

It is against the rules to give a direct answer but I always conffirm an answer. I dont see wht's wrong with that if the questioner does some calculations themself.